OpenStudy (anonymous):

Write binary number \[10,\overline{10111}\], as decimal fraction and reason-show it via geometric series

OpenStudy (anonymous):

is the , meant as a decimal?

OpenStudy (anonymous):

hmm as decimal fraction i tink something like \[\frac{x..}{x..}\]

OpenStudy (anonymous):

\[x=\{0,1,2,3,4,5,6,7,8,9\}\]

OpenStudy (anonymous):

no, what I meant is \[10. \overline{10111}=10, \overline{10111}\]

OpenStudy (anonymous):

yes you are right i think its both ok

OpenStudy (anonymous):

in question table given as with comma ","

OpenStudy (anonymous):

so 10 = 4, yeah? \(.\overline{10111}=23/32\)?

OpenStudy (anonymous):

in binary i think \[10 = 2^{1}+{2}^{0} = 2 + 1 = 3 \]

OpenStudy (anonymous):

and as such, \(.\overline{0000010111} = \Large{\frac{23} {32^2}}\)

OpenStudy (anonymous):

.10111 = 1/2 + 0/4 + 1/8 + 1/16 + 1/32 = (16 + 0 + 4 + 2 + 1)/32 = 23/32

OpenStudy (anonymous):

in binary, \(.1 = 1 \times 2^{-1}\text{, }.01 = 1 \times 2^{-2}\text{, }.001 = 1 \times 2^{-3}\text{, and so on}\)

OpenStudy (anonymous):

and \[\frac{1}{2}+\frac{0}{4}+... \] is the geometric serie right ?

OpenStudy (anonymous):

no... \(.\overline{10111} = .101111011110111...\) what repeats is blocks of 5, which in binary is powers of 32. so you have to break the fractional part (the part after the decimal sign) into groups of 5. \[.10110=\frac{23}{32} \text{, }\]\[.0000010111=\frac{23}{32^2}\text{, }\]\[.000000000010111=\frac{23}{32^3}\text{, and so on...}\]

OpenStudy (anonymous):

ok pgpilot thank you very much

OpenStudy (anonymous):

does that make sense? do you see what the geometric part is?

OpenStudy (anonymous):

geometric series i think \[32^{1}, 32^{2}, 32^{3}\] right ?

OpenStudy (anonymous):

no, because the powers of 32 are in the denominator...

OpenStudy (anonymous):

a ok like \[(2^{4})^1,(2^{4})^2 ... \] and so on right ?

OpenStudy (anonymous):

sorry \[(2^{5}).... \]

OpenStudy (anonymous):

\[4 + \frac{23}{32} +\frac{23}{32^2}+\frac{23}{32^3}+\frac{23}{32^4}+ \cdots\]

OpenStudy (anonymous):

now its more clear to understand thank you very very much pgpilot